OpenStudy (anonymous):

Identify the 17th term of a geometric sequence where a1 = 16 and a5 = 150.06. Round the common ratio and 17th term to the nearest hundredth. someone answer please

OpenStudy (anonymous):

the n-th term of a geometric sequence is\[a_n = a_1 r^{n-1}\]use that equation to find the common ratio (r) and the 17th term (a17)

OpenStudy (anonymous):

im getting stuck at solving for r

OpenStudy (amistre64):

that parts just algebra:\[\frac{a_n}{a_1}=r^{n-1}\] \[\large \sqrt[{n-1}]{\frac{a_n}{a_1}}=r\]

OpenStudy (anonymous):

4 \[\sqrt{150.06}\] /16 = r ?

OpenStudy (amistre64):

yes, if i read that right :) lets use exponents instead :)\[\sqrt[a]{b}=b^{1/a}\] \[(\frac{150.06}{16})^{1/4}\]

OpenStudy (anonymous):

whenever i divide 150.06 by 16, i get 9.37875 or 9,.38 i just dont know how to ger rid of the 4 in front of the square root sign

OpenStudy (anonymous):

wait how do you get rid of the 1/4 ?

OpenStudy (amistre64):

you dont get rid of it, you use it

OpenStudy (anonymous):

well how do you use a fraction exponent ?

OpenStudy (amistre64):

are you doing this by hand, calculator, or website assisted?

OpenStudy (anonymous):

hand and calculator

OpenStudy (amistre64):

by hand is almost impossible do you know where the exponent buttom on your calc is? should look like "^"

OpenStudy (amistre64):

other buttons you need are "(" "/" and ")"

OpenStudy (anonymous):

when i plug it in to mu calculator i get 2.3446875

OpenStudy (anonymous):

or 2.34

OpenStudy (amistre64):

ill use google to verify :) http://www.google.com/search?rlz=1C1CHFX_enUS482US482&sourceid=chrome&ie=UTF-8&q=(150.06%2F16)%5E(1%2F4) 1.74999...

OpenStudy (anonymous):

so r is 1.74999 ? or 1.75 ?

OpenStudy (amistre64):

round as it tells you to round, but yes

OpenStudy (anonymous):

so then i plug back into orignial equation ?

OpenStudy (amistre64):

well, that is why we wanted it to begin with, so that we could use it. so im going to go with a yes

OpenStudy (anonymous):

okayy

OpenStudy (amistre64):

a17 = 16 r^16 if you need a little guidance ;)

OpenStudy (anonymous):

i got the answer thank you